Find the domain of x-2y=1 if the range is (-2,0,5,8,)
Answer by cngriffith(27)

The domain is the x value. The range is the y value. Alphabeticly domain comes before range. Alphabaticly, x comes before y. Therefore x=domain and y= range.
x-2y=1
x=2y+1 Solve equation for x.
x=2(-2) +1
x=-4+1=-3 Substitute the first value in the range for y and solve for x.
Continue this process for each number in the range.
x=2(0)+1
x=1
x=2(5)+1
x=11
x=2(8)=1
x=17
Domain is -3,1,11,17